### 05 Feb 2024 v1

# ReviewFlow: Intelligent Scaffolding to Support Academic Peer Reviewing

Abstract.
Peer review is a cornerstone of science. Research communities conduct peer reviews to assess contributions and to improve the overall quality of science work. Every year, new community members are recruited as peer reviewers for the first time. How could technology help novices adhere to their community’s practices and standards for peer reviewing? To better understand peer review practices and challenges, we conducted a formative study with 10 novices and 10 experts. We found that many experts adopt a workflow of annotating, note-taking, and synthesizing notes into well-justified reviews that align with community standards.Novices lack timely guidance on how to read and assess submissions and how to structure paper reviews. To support the peer review process, we developed ReviewFlow – an AI-driven workflow that scaffolds novices with contextual reflections to critique and annotate submissions, in-situ knowledge support to assess novelty, and notes-to-outline synthesis to help align peer reviews with community expectations. In a within-subjects experiment, 16 inexperienced reviewers wrote reviews using ReviewFlow and a baseline environment with minimal guidance. Participants produced more comprehensive reviews using ReviewFlow than the baseline, calling out more pros and cons, but they still struggled to provide actionable suggestions to address the weaknesses. While participants appreciated the streamlined process support from ReviewFlow, they also expressed concerns about using AI as part of the scientific review process. We discuss the implications of using AI to scaffold peer review process on scientific work and beyond.
intelligent scaffolding, academic peer review, Large Language Models LLMs)
copyright: acmcopyrightjournalyear: 2024price: 15.00ccs: Human-centered computing Empirical studies in HCI
1. Introduction
Peer review is a cornerstone of academic research, ensuring the quality, credibility, and reliability of scientific research (Price and Flach, 2017). The peer review process seeks to assess whether submissions contribute new knowledge to a research community and generate feedback that helps authors improve the quality of their work  (Jefferson et al., 2002; Shah, 2022). Many communities are seeing a rapid increase in submissions (McCook, 2006; Stelmakh et al., 2019; Shah, 2022; Arous et al., 2021; Stelmakh et al., 2021a); while this could be seen as an indicator of scientific progress, it also has increased the pressure on peer review ecosystems. To meet increased demand, many research communities recruit a significant number of first-time reviewers for each review cycle. For example, in 2023, the ACM CHI Conference on Human Factors in Computing Systems reported that more than 50% of ACs who reviewed papers are first-time ACs (chi, 2015).
Peer reviewing is a complicated and challenging task. Reviewers need to understand the paper, evaluate the scientific content using domain knowledge, make a fair decision, and compose a comprehensive draft to communicate their assessments and recommendations (Shah, 2022). It is a task that requires a deep understanding of the subject, the critical thinking process, and the ability to provide constructive feedback. Conferences and journals need to ensure that first-time reviewers meet the standards of the research community.
Figure 1. ReviewFlow interface. Users can (A) receive section level reflection cues guided by community criteria; (B) request phase level reflection cues adapted to the highlight paper content; (C) click the citation to check in-situ summarization; (D) check the recommended citations that are potentially missed by the paper; (E) click to summarize the notes into high level structured outline and expand it to a detailed outline
One approach to preparing novices for any complex task is to provide scaffolding, a strategy used in educational settings to support learning and mastery (Reiser, 2004; Collins, 2006). Scaffolding takes many forms, including examples of prior work, templates, or hints to help novices think more and potentially perform more on par with experts (Holton and Clarke, 2006; Saye and Brush, 2002). Previous studies found that scaffolded examples and templates can even help learners perform similarly to experts in terms of feedback quality (Yuan et al., 2016). Recently, LetterSmith’s instructional approach to aid writing called “scaffolded annotation” which provided key components of the writing tasks and annotated expert examples, helped professional writing students improve the quality of their early-stage drafting in a new genre (Hui and Sprouse, 2023). Advances in AI and large-language models (LLMs) in particular, have the potential to make scaffolding even more effective, because they can interpret and adapt to the user’s work context (Rachatasumrit et al., 2022). For example, the CReBot project leverages LLMs to generate guiding questions to help induce creative thinking while reading academic papers (Peng et al., 2022). Experimental results showed that CReBot can engage novices in reading and help them comprehend the paper content. While LLMs show great potential, they are also known to hallucinate (Bang et al., 2023) and perpetuate systemic biases (Huang et al., 2019; Jakesch et al., 2023) that could negatively impact the effectiveness of scaffolding. While intelligent scaffolding has proven valuable for well-scoped tasks, academic peer review involves reading submissions with a critical eye, synthesizing knowledge, and making a well-justified judgment for acceptance or rejection. Our research investigates how intelligent scaffolds can guide a complex, multi-faceted workflow for academic peer reviewing without biasing the ultimate decisions. This paper explores two key research questions:
  • What are the strategies adopted by experts and challenges faced by novices during the peer review process?
  • How can we leverage LLMs to intelligently scaffold the peer review process for novices without introducing bias or over-reliance?
To explore how we might use AI scaffolding to support the peer review process, we first conducted a formative study to understand what novices see as key challenges and how experts approach this task. We interviewed 10 novices with limited prior experience writing peer reviews to articulate their obstacles during the process. They expressed challenges around the lack of sufficient guidance on how to write a well-structured peer review peer review and how to make judgments about the paper’s quality. Furthermore, we conducted observational studies with 10 experts to ask them to write a peer review for a short paper and to think aloud as they complete their workflow. Then, we invited experts to provide their perspectives on using AI to support the tasks and express their concerns. We found that many experts adopt a workflow involving critical reading, annotating, note-taking, and synthesizing a well-justified review that conforms to community guidelines.
Based on these insights, we developed a prototype – “ReviewFlow” [^1] ReviewFlow Code Repository:https://github.com/creativecolab/ReviewFlow.git, a platform for writing peer reviews that incorporate intelligent scaffolding to support a workflow for inexperienced reviewers. ReviewFlow incorporates a range of features to facilitate the review process: (1) Contextual cues are embedded questions that help reviewers reflect on the paper. ReviewFlow includes section-level cues guided by the community’s review criteria as well as phrase-level cues that adapt to the paper’s content. (2) In-situ citation recommendations show relevant but non-cited papers that may help reviews assess the novelty compared to existing work. (3) Notes-to-outline synthesis guides reviewers to organize notes and structure reviews to align with community standards. ReviewFlow gathers all the notes left by the reviewer and leverages an LLM to summarize notes into a high-level outline. Reviewers can revise and add detail to the outline while adhering to community standards.
We conducted a within-subjects study to evaluate ReviewFlow where (N=16) participants — with little to no experience as reviewers — wrote reviews for two short papers in a counterbalanced manner: one using the ReviewFlow with full scaffolding features and one using the baseline interface that only has the traditional forms of guidance(e.g., review rubrics and an example review). We found that novice reviewers wrote significantly more structured and more comprehensive reviews in the ReviewFlow system than in the Baseline system, evaluated by experts. Novice reviewers wrote slightly more constructive reviews in the ReviewFlow system, but the difference was not significant. Reviewers called out more weaknesses in the paper using ReviewFlow, but they still struggled to provide constructive and actionable suggestions for the authors to overcome the weaknesses.
Our paper offers several contributions: First, a formative study revealed that novices lack opportune guidance on key considerations and expectations and uncovered common practices adopted by experts in the review process. Second, we developed ReviewFlow to model experts’ workflow for peer reviewing while also leveraging LLMs to provide contextual cues, in-situ knowledge recommendations, and notes-to-outline synthesis. Third, we gained empirical insights from a within-subjects study with 16 participants which revealed how intelligent scaffolding can help novices write well-structured and comprehensive reviews.
2. Related Work
2.1. Practices and challenges related to academic peer reviewing
Peer review is an essential step of academic research. As an important step for ensuring the scientific quality of work, peer review has been adopted by most journals and conferences (Price and Flach, 2017). In a typical conference review process, each reviewer needs to evaluate the paper’s quality, make an acceptance decision and provide reviews for their assigned papers (Shah et al., 2018; Shah, 2022). The reviewers are usually experts in the area who have fruitful experience and knowledge to assess or evaluate the quality and contribution of the paper. After reviewers submit the review, a discussion takes place between reviewers and a meta-reviewer, who will carry out the final decision on the acceptance of the paper. Note that conferences may have some differences in their peer-review process.
The number of papers in research communities has increased exponentially in recent years (Wang et al., 2020). While this may be positively viewed as an acceleration of scientific progress, the disparity between growth rates of the submission and reviewer pools also creates more burden for reviewers (McCook, 2006; Stelmakh et al., 2019, 2021a; Shah, 2022). To avoid overloading reviewers, conferences need to find new sources of reviewers as there are not enough experienced reviewers to review all papers (Stelmakh et al., 2021a). These novice and junior reviewers constitute a large fraction of the reviewer pool in computer science conferences (Stelmakh et al., 2021b). For example, using data on the structure of the reviewer pool of the ICML 2020 conference, researchers estimated that approximately 35% of the reviewers were novice reviewers who are junior individuals who self-nominated and satisfied the screening requirements of having one or two papers published in some top ML venues (Stelmakh et al., 2021b). Given this large fraction, conferences need to ensure that newly added junior reviewers do not compromise the quality of the process, that is, are able to write reviews of quality comparable to the experienced reviewers.
A previous study explored and compared the reviews written by experienced reviewers versus junior reviewers and the study showed that junior reviewers were slightly harsher in scoring the clarity of the submissions (Shah et al., 2018; Stelmakh et al., 2021a). In the meantime, other works provide empirical evidence that junior reviewers are more critical than their senior counterparts and reveal that graduate students’ review comments are not very useful (Mogul, 2013; Patat et al., 2019). Faced with these doubts and challenges, helping novice junior reviewers to write a high-quality review becomes crucial.
Reviewing is a complex, time-consuming, and mentally demanding task (Wang et al., 2020; Shah, 2022). A constructive and comprehensive review can improve the quality of the paper, while a bad, random, dismissive, or biased review brings frustration and anger to reviewers (Wang et al., 2020). To provide a high-quality review, reviewers need to go through a multi-step workflow: first, they need to read and fully understand the contribution of the paper; second, they need to use the domain knowledge to assess the merit of scientific contributions and make a decision on the submission; last, they need to provide an evaluation that covers multiple aspects, including the originality, clarity, validity, etc (Smith, 2006; Jefferson et al., 2002; Xiong and Litman, 2011; Langford and Guzdial, 2015; Hua et al., 2019; Cheng et al., 2020; Gao et al., 2019; Wang et al., 2020; Yuan et al., 2021). To make a fair judgment, reviewers need to have enough background knowledge to grasp the main idea of the paper and evaluate its contribution. More importantly, reviewers need to equip critical thinking skills to think deeply about the author’s judgments, like whether the claims are reasonable and why the approaches are chosen (Facione, 1990; Moore, 2013). A typical peer review not only contains the paper summary and its contribution but also raises weaknesses from different aspects together with constructive feedback or thought-provoking questions. During this process, it requires readers to actively analyze, synthesize, and evaluate the paper content (Moore, 2013).
Researchers are mainly trained to do research while there is little training for the peer review process. Hence, it becomes even more challenging for inexperienced junior reviewers to gain expertise quickly (Shah, 2022). Existing research explored instructional methods to “teach” or “train” junior reviewers. A previous study provided a training video and found that it increased the inter-reviewer agreement, alignment with the scoring rubric, and the amount of time reading the review criteria (Sattler et al., 2015). Another study offers novice reviewers a more guided introduction to the different stages of the reviewing process, such as how to lead a discussion among reviewers, with the goal of helping novices to write better reviews. The results showed that with this guidance on the reviewing stages, novice reviewers can deliver more “above expectation” reviews (Stelmakh et al., 2021a). However, their guidance is only limited to introducing the different parts of the reviewing process, such as rebuttal and discussion, and providing novices opportunities to ask expert questions on the general process, for instance, how to initiate a discussion among reviewers (Stelmakh et al., 2021a). Outside of the general review process, academic peer review is a complex activity that involves both understanding a paper submission’s stated contributions and evaluating whether the paper crosses an acceptable threshold for the research community. To explore how we might support the peer review process, our research starts with a formative study to understand how experts approach this task and what novices see as key challenges.
Previous research has used computational methods to provide support to streamline several parts of the peer review process, such as matching submissions with appropriate peer reviewers, authoring more comprehensive and decisive reviews, as well as review quality assessment (Bartoli et al., 2016; Wang et al., 2020; Hua et al., 2019; Yuan et al., 2021; Arous et al., 2021; Shah et al., 2018). However, there is less empirical study that attempts to scaffold the entire workflow for reviewing academic papers, which includes reading, note-taking, evaluating, decision making, and synthesizing this into a written review.
2.2. Scaffolding strategies for complex cognitive tasks
To help novices improve problem-solving skills in complex cognitive tasks, cognitive apprenticeship introduces several strategies, including modeling, coaching, scaffolding, and reflection (Vygotsky and Cole, 1978; Collins, 2006). Scaffolding is instructional support provided by experts to promote learning, especially when concepts and skills are being first introduced to novice students (Reiser, 2004; Collins, 2006). These supports include advanced organizers, modeling, worked examples, concept maps, explanations, handouts and prompts (Oakes et al., 2018; Nesbit and Adesope, 2013; Atkinson et al., 2000; Charney and Carlson, 1995; Bui and McDaniel, 2015; Collins, 2006). Previous research shows that effective scaffolding can help novices perform work nearly as well as experts (Yuan et al., 2016; Hui and Sprouse, 2023; Hui et al., 2018). When scaffolding is mediated by technology, including AI-based methods, it creates more opportunities for instructors and learners but also brings more challenges in making the scaffolding contextualized, adaptive, and effective (Hannafin et al., 1999).
Researchers used prompts and guided questions to scaffold learners in the paper reading process (Peng et al., 2022; Yuan et al., 2023; Chen et al., 2022). To facilitate critical paper reading, researchers developed CReBot which interactively asks section-level critical thinking questions for routine paper readers and novice readers. Results showed that the interactive question prompts CReBot provided might not be better than static guidelines for beginners to read. Similarly, researchers developed CriTrainer which can adaptively provide questions in the reading process together with hints and feedback to help readers critically think and comprehend the paper content. Interestingly, on the opposite of CReBot, their result showed that CriTrainer can improve learners’ ability to raise understandable, relevant, and critical questions after the training sessions. Their results highlighted the benefits of its text-specific critical thinking questions provided by the system. However, guided questions used in CReBot and CriTrainer are both template-based, which ignored the contextual information provided by the paper.
Existing research also uses existing examples and templates to scaffold complex writing process (Yuan et al., 2016; Hui and Sprouse, 2023; Hui et al., 2018). In the writing scenario, scholars found that scaffolding can help students learn about form and organization from analyzing the examples and templates (Doan, 2021; Collins, 2006). In the context of writing introductory help requests, providing high-quality examples and expert informed templates can increase learning and writing quality (Hui et al., 2018). Another writing support system used scaffolded annotation and examples as an instruction approach to professional writing students improve the quality of their early-stage drafting in a new genre (Hui and Sprouse, 2023).
In traditional instruction scenarios, experts take a huge amount of time to curate examples, create guidance, or author rubrics (Reiser, 2004). However, experts who created examples still faced the challenges of effectively adapting and contextualizing into the current learning step. Thus, this brings AI and machine intelligence opportunities to provide efficient and context-specific scaffolding in the learning process (Wambsganss et al., 2020, 2021). Previous research has developed multiple methods for using machine intelligence to scaffold complex tasks, including writing and researching (Norkute et al., 2021; Rachatasumrit et al., 2021; Yuan et al., 2022; Gero et al., 2021). For example, researchers developed ArgueTutor, a conversational agent that tutors students with adaptive argumentation feedback in their learning journey using Natural Language Processing (Wambsganss et al., 2021). However, as far as we know, none of the studies that used scaffolding strategies focused on the context of conference peer review writing. Our research explores how we might guide a complex, multi-faceted workflow like peer reviewing and the potential role of AI in creating contextually adaptive scaffolds.
2.3. Leveraging AI to support writing
Recent works explored the potential of using AI to support people’s workflow on writing tasks (Ippolito et al., 2022; Chung et al., 2022; Clark et al., 2018; Gero and Chilton, 2019b, a; Rahman et al., 2020; Lee et al., 2022; Gero et al., 2021; Ippolito et al., 2022; Mirowski et al., 2022). For example, TaleBrush allows users to create a story with AI through sketching to aid the planning of writing. Then it enables writers to generate diverse storylines and interactively refine them (Chung et al., 2022). Another system Wordcraft explores how to support users collaborating with generative language models to co-write a story (Yuan et al., 2022). Spark used a language model to generate prompts related to a scientific concept to facilitate scientific writing (Gero et al., 2021).
Research showed that writing is a complex, iterative process (Flower and Hayes, 1981). The Hayes model describes the cognitive processes an individual writer engages in during the process of writing (Flower and Hayes, 1981; Hayes, 2012; Gero et al., 2022). In the writing process, there are three major components: the planning component, the translating component, and the reviewing component, as shown in Figure 5. To facilitate the cognitive process of writing, several systems are developed to support different stages (Bhat et al., 2023; Zhang et al., 2023). For example, VISAR is an AI-enabled writing assistant to helps writers brainstorm and revise hierarchical goals and organize argument structures in the planning stage.
To facilitate the iterative planning and revising process in writing, intelligent systems further use the chain of thoughts (COT) prompting method to break down the large problem into step-by-step prompts (Wei et al., 2022). Specifically, Re3 framework and DOC framework used the COT approach to decompose the writing tasks where they first generate an outline and then automatically turn the outline into the story generation. The evaluation demonstrated that this decomposition approach can improve the coherence of long story generation (Yang et al., 2022b, a). However, this approach could be highly controllable where humans can control the story generation by modifying the outlines. Control and agency are extremely important in the conference peer review writing where human reviewers should play the role of driving the writing process. The development of LLM brings opportunities along with concerns on what LLMs shouldn’t do in the academic writing area (Fok and Weld, 2023). The prior work that aligns most closely with the concept of applying AI techniques to reviewing academic papers that a machine model that automatically generates feedback using LLMs (Liang et al., 2023). Results showed that LLM feedback could benefit researchers in earlier stages of manuscript preparation while researchers struggle with an in-depth critique of study methods. Instead of using an automatic method, humans should be in the loop to drive and control the writing process (Fok and Weld, 2023). Drawing on these insights, we designed the ReviewFlow system not to automate any parts of the process, but rather to scaffold key considerations and to give novices agency over how to apply machine-generated language suggestions.
3. Formative Studies
Before we developed our system to support academic peer reviewing, we conducted a series of formative studies. First, we interviewed ten novice reviewers to understand how they approached this task for the first time and their perceived challenges(represented as C). Second, we observed ten experienced reviewers as they conducted a peer review on a couple of selected papers to understand common practices and workflows. Last, we propose design goals((represented as DG)) on supporting novice reviewers.
3.1. Methods
3.1.1. Novice Interview Study
We conducted 30-minute semi-structured interviews with 10 novice reviewers with relatively little experience with writing academic peer reviews (ranging from only reviewing once before and having less than 2 years of experience). Participants were recruited through mailing lists and social media posts. The participants (4 female and 6 male, average age of 25.5 years) came from diverse research fields including Human-Computer Interaction, Artificial Intelligence, Cognitive Science, and Computer Science. Interviews were conducted remotely by the lead author and lasted around 30 minutes.
In the interview, we first asked open-ended questions about their perceived obstacles and challenges of conducting academic peer review. We then provided scenarios of potential features for peer review scaffolding to elicit reactions from novice reviewers. These scenarios described common potential situations faced by first-time reviewers and used these to elicit participants’ needs in a real-world situation (Gaver et al., 1999). For example, “Mary is a first-year graduate student doing Human-Computer Interaction (HCI) research. While she has submitted a couple of papers before she received reviews, this will be her first time as a reviewer. She struggled to make the review constructive for the authors.”. A complete list of scenarios and design probes is listed in the Appendix. Based on these answers, three researchers on our team iterated multiple times to come up with design probes targeting these challenges and visualize them as design probes in Figma.
In the end, we provided a series of design probes for the novices to capture their reactions. Drawing insights from previous literature around the intelligent support on reading and writing, the research team took multiple rounds of iteration and group discussions to develop the six design probes (Zhang et al., 2023; Lo et al., 2023). The design probes included: (1) scaffold annotation with community curated tags and filters, (2) contextual reflection cues (expert-authored generic cues versus AI-generated specific reflection cues on each section or highlighted sentences), (3) extractive summarization and generated explanations to facilitate paper reading, (4) in-situ citation recommendation where the system provides a summary from the cited paper and recommend potential missing citations, (5) review draft generation with structures, (6) mapping back the source of review draft from the paper content and visualize the location of the source to help reviewers revise their draft. Interviews were recorded with participants’ permission and were transcribed into text. Three researchers from the team went through the transcripts and coded them for themes using thematic analysis (Braun and Clarke, 2006). Through multiple iterations along with periodic discussions with the rest of the research team, the coding led to the major themes of challenges below.
3.1.2. Experts Observational Study
We also chose to recruit experienced peer reviewers to learn about their best practices through a contextual inquiry (Karen and Sandra, 2017). We recruited 10 experts with at least 5 years of experience in writing academic peer reviews. Participants were recruited through in-person invitations, email lists, and social media posts. Participants (4 female and 6 male, average age of 30.1 years) came from different fields of computer science, including Human-Computer Interaction, Artificial Intelligence, programming languages, learning science, computer security, and accessibility. Observational studies were conducted remotely by the lead author and lasted around 90 minutes.
During the observational study, we first spent 20 minutes asking about their experiences and challenges with peer reviewing. Then, the team asked the participant to write a peer review in a Google document for one of 5 different short papers (less than 4 pages) in 60 minutes. To find suitable short papers, the team first collected the participants’ research interest descriptions from their websites and used these keywords to search on Semantic Scholar (sem, 2015). We filtered out the papers longer than 5 pages, ranked them by relevance, and selected the top N as the participant’s options for the peer review exercise. Last, we provided the same design probes to the experts to elicit what scaffolding strategies could benefit the novices the most. The interviews were recorded and then transcribed using a machine transcription service. The research team took the same analysis procedure as the previous novice study.
3.2. Findings
3.2.1. Novice reviewers felt they lacked guidance on evaluating the paper and writing structured and constructive reviews
The novice participants reported that their prior attempts at reviewing papers were cognitively demanding and took an average of 6.4 hours. When asked about challenges, 6/10 mentioned that they struggle to fully understand the background knowledge and existing work and did not feel confident about assessing the paper’s novelty [C2]. 4/10 mentioned that they lacks opportune considerations for assessment [C1]. Specifically, one participant mentioned that they would love to have “co-reviewers who have already read the paper and known the specifics to guide me through the evaluation process”(N5). In addition, 4/10 highlighted that they need more guidance on how to structure issues during the writing process[C3]. 3/10 explicitly mentioned that they want to receive some feedback from experts, other reviewers, or authors, to make sure the review is in the right tone and meet the expectation [C4]. Specifically, one participant mentioned “I am not sure whether I covered all the necessary points or whether authors will perceive the reviews as useful”(N7). Another participant is worried about being “impolite or harsh (N9)”.
We presented novices the design probes (see Appdendix) one by one and asked them to rate the degree to which they resonate with each scenario (1=Not resonate at all to 5= Strongly resonate with the situation). The top two situations that resonate with novice reviewers the most were: “the novice reviewer struggles to write a high-quality review”(3.8/5) and “the novice reviewer spends a lot of time reading a paper but doesn’t know how to evaluate the paper”(3.8/5).
3.2.2. Experienced reviewers adopt a workflow of sense-making, annotating, and synthesizing notes
Experienced reviewers reported that they spent 4.75 hours to review one paper. In the observational study with 10 experts, we synthesized the common workflow adapted by experts in the reviewing process. As an initial step, reviewers read and comprehend the paper’s content. While reading through the paper, all experts (10/10) highlighted some content, annotated sentences or paragraphs, and took some notes. The format and style of notes varied between different reviewers. We observed some reviewers using symbols while others used phrases or short sentences. After reading and annotating the content, reviewers start to re-read the notes to evaluate the paper’s quality. 3 out of 10 experts went back to check the introduction and related work to assess the novelty of the paper.
After finishing reading the paper, instead of directly editing the review draft, experts created high-level headers or topics, such as “lack clarity on study design”, that summarized the paper’s weaknesses and strengths as well as prioritized the concerns (7/10). For instance, one expert reflected on the review process as “I will make a lot of annotations and then I will just review them section by section. At the same time, thinking about what is the biggest concern in terms of the overall research novelty”(E6). We observed that 6 out of 7 participants wrote the review in the following structure: summary of the paper, strengths or contribution of the paper, two to three weaknesses of the paper, and end the review with decision justification and general recommendations. Some experts will then list bullet points under each topic together with questions that they would like to ask the author. Last, experts compiled these comments and bullet points into a complete draft and revised them two to three times to offer suggestions and make it more constructive. Figure 2 represents the expert’s review workflow together with their practices.
3.2.3. Experienced reviewers stressed the importance of specific and contextual guidance for scaffolding
Among these features, 5 experts reflected that the contextual cues can be helpful, especially for novice reviewers. One participant preferred the AI-generated cues on the selected paper content and explained that “I like to have some capability of freedom, but I think this will be super useful for novice reviewers who don’t know how to review. But for people who have reviewed for so many years, those guidances for conferences are pretty much the same”(E6). We provided experts with two sets of tags to select. The first set of tags is designed based on the review structure that includes “summary of the paper”, “strength”, “weakness” and “others”. The second set of tags is designed using community review criteria that include “relevance”, “novelty”, “validity”, “clarity”. Most experts preferred the first set of tags (7/10). 2 participants mentioned that they were concerned the experts’ authored cues might be too similar to the existing guidelines. Hence, they suggested the contextual question can provide better guidance for novices.
9 out of 10 experts mentioned their preference for the in-situ citation support to provide summary and recommendations. They reflected that this in-situ knowledge support can “raise awareness on unknown work”(E5). 9 out of 10 participants expressed their concerns about the summarization feature where they don’t trust the AI’s ability to identify important information. Instead, they think reviewers should have control over the reading process. Specifically, one expert mentioned “I think we are also reviewing the style of writing, or how something is communicated and logically connected between each paragraph or each sentence. So I think there is value with actually reading everything to get the message behind the paragraphs”(E3). Participants shared their opinions on using AI in the review process and all of them agreed that LLM should not generate the review on the fly while human experts should drive the process since the limitation of LLM can bias human experts and bring over-reliance on the use of AI.
Based on the interview on the design probes, we summarize the following design considerations mentioned by novices and experts. The support should be lightweight and don’t distract the current review flow. While involving AI in the review process, AI should not go too far to biases or lead the thinking process. Human reviewers should still preserve agency in reading, writing, and decision-making. Faced with the limitations of current LLM, intelligent systems should try to avoid hallucinations and provide enough opportunities for fact-checking.
3.3. Formative Study Discussion
From the interview study with novices and the observational study with experts, we identified a range of challenges perceived by novices as well as insights on expert practices (see Figure 2). Novices lack confidence in identifying the novelty of research based on prior research and in knowing how to structure a peer review that meets community standards. Experienced reviewers tend to avoid biasing the decision-making and preserve agency in reviewing. The juxtaposition of these novice challenges and the expert practices suggests four core design goals for intelligent scaffolding:
Figure 2. Experts’ workflow in academic peer review, along with experts’ practices, novices challenges, and design goals for each stage
DG1: Offer lightweight and adaptive scaffolds that facilitate reflection throughout the paper Novices highlighted their need for guidance in critically evaluating papers from various perspectives, expressing uncertainty about the key points to focus on [C1]. In educational settings, instructors typically serve as scaffolds when introducing new concepts or knowledge to novice learners (Collins, 2006). Prior studies in scientific paper reading developed methods to  (August et al., 2022; Rachatasumrit et al., 2022). For example, Paper Plain provided a collection of key questions that guide readers to answering passages and plain language summaries of those passages (August et al., 2022). Similarly, in the context of our study, the objective is to provide guidance comparable to that of expert peer reviewers. The emphasis is on delivering locally relevant questions that help users think critically about each section of the paper.
DG2: Enable in-situ knowledge support for assessing novelty compared to prior work Novices are struggled with insufficient background knowledge for evaluating papers [C2]. Providing knowledge support in situ can externalize the user’s working memory, aid in sense-making and facilitate a swift reviewing and resumption of task contexts (Kuznetsov et al., 2022). Prior studies in scientific literature review highlight the importance of in-situ knowledge support (Kang et al., 2022; Chang et al., 2023). To help novice reviewers who lack background knowledge to evaluate the novelty while reading the paper, our goal is to enable in-situ knowledge support by extracting the abstract of the cited paper.
DG3: Model the expert workflow of reading, evaluating and synthesizing practiced by experienced reviewers Novices expressed their need for more specific guidance from experts in terms of their process and structure [C3]. Prior work showed that surfaced expert practices can better structure and scaffold the process for novices (Kim et al., 2015). Motify illustrates that storytelling patterns extracted from expert stories can be used to effectively scaffold novices to create video stories (Kim et al., 2015). In the context of writing introductory help requests, providing high-quality examples and expert-informed templates can increase learning and writing quality (Hui et al., 2018). Our approach involves the modeling of the expert workflow encompassing reading, evaluating, synthesizing, and revising. This model is then employed to structure the peer review process for novice reviewers.
DG4: Guide the reviewing process such that the work aligns with community standards Novices reported that they lacked effective models for tone and expectation [C4]. More guidance on the expectation of the review can help reviewers to reflect on their tone and structure. In addition, experts raised valid concerns about the potential for intelligent tools to introduce biases or influence the thinking process. Existing studies have highlighted apprehensions regarding AI exhibiting biases in human-AI collaboration tasks (Kawakami et al., 2022; Poursabzi-Sangdeh et al., 2021). Notably, research has shown that providing information, including explanations, generated by AI has the potential to mislead users in decision-making (Lakkaraju and Bastani, 2020; Goyal et al., 2023; Buçinca et al., 2021). Therefore, a key design objective is to steer clear of biasing the decision-making process and focus solely on encouraging justifications. This approach aims to assist novice reviewers in producing high-quality and original reviews without compromising the integrity of the evaluation process.
4. ReviewFlow
We developed ReviewFlow which employed AI-driven scaffolding strategies naturally into the review workflow to support novice reviewers to gain expertise in conference peer reviewing.
4.1. Key features
4.1.1. DG1: Contextual cues
Researchers used prompts and guided questions to scaffold learners in the paper reading process (Peng et al., 2022; Yuan et al., 2023; Chen et al., 2022). Understanding the paper and critically reflecting on the content are essential for decision-making and review writing in the later stage. Previous research showed that providing questions can increase user engagement in actively searching for information. Previous systems mostly use template-based guided questions to engage readers (Peng et al., 2022). To Offer lightweight and adaptive scaffolds that facilitate reflection throughout the paper [DG1], ReviewFlow provides two types of contextual cues for novices, as shown in Figure 3. One type is section level reflection cues guided by community criteria which are adapted based on each paper section’s content together with the review criteria. For section-level reflection cues in the Figure 7-A), it took the entire section, the paper abstract, and the community criteria into account to generate questions. Readers can either read these questions before they read the section to obtain contextual guidance or after to reflect on the section’s content. Another type is phrase level cues adapted to paper content. As shown in Figure 7-B, participants have the freedom to highlight the content that they would like to reflect on at the phrase level and then select the review criteria. Then, it generated cues for participants to reflect on.
Figure 3. Contextual cues. ReviewFlow provides (A) section level cues guided by community criteria and (B)phrase level cues adapted to paper content and selected criteria
4.1.2. DG2: In-situ citation recommendation as knowledge scaffolding
Previous research in scientific literature reviewing highlights the importance of in-situ knowledge support (August et al., 2022; Kang et al., 2022; Rachatasumrit et al., 2022; Chang et al., 2023). For example, PaperPlain provided in-situ plain language section summaries to facilitate paper reading. To enable in-situ knowledge support for assessing novelty compared to prior work, ReviewFlow in-situ presents TLDR summary when the user click on each reference to provide background knowledge around each paper, as shown in Figure 4-D. The TLDR is queried from semantic scholar (sem, 2015). To further raise the awareness of the existing reference in the reading workflow, ReviewFlow provided a popup window with potential missing citations from the same venues, as shown in Figure 4-C.
Figure 4. In-situ knowledge scaffolding. ReviewFlow provides: (C) a list of relevant paper from the same venue that are not cited and (D) in-situ citation summary including title, author and TLDR
4.1.3. DG3: Notes-to-outline synthesis
In the review process, we observe expert reviewers synthesize notes to make a plan on how to draft a review. To facilitate the iterative planning and revising process in writing, intelligent systems further use the chain of thoughts prompting method to break down the large problem into step-by-step prompts (Wei et al., 2022). Specifically, previous research used the chain of thought approach to decompose the writing tasks where they first generate an outline and then automatically turn the outline into the story generation. The evaluation demonstrated that this decomposition approach can improve the coherence of long story generation (Yang et al., 2022b, a). However, this approach could be highly human controllable where humans can control the story generation by modifying the outlines. Recent studies explored the potential of automatically generating feedback using LLM for conference papers (Liang et al., 2023).
Figure 5. Scaffolding features support the review drafting process. The top part shows the Flower and Hayes cognitive process of writing (Flower and Hayes, 1981). The bottom part shows the ReviewFlow features that support each stage of writing. On the left, ReviweFlow summarized notes into broad topics under strengths and weaknesses to help reviewers plan the writing. In the middle, the topics can expand to a detailed outline. On the right, it encourages self-reflection and allows post-editing
By modeling the expert workflow of reading, note-taking, and synthesizing practiced by experienced reviewers, we designed the notes-to-outline synthesis feature that facilitates notes synthesis and review planing (Flower and Hayes, 1981). Figure 5 showed that ReviewFlow first synthesized notes to high-level outline to help reviewers organize their notes into structured topics and provide ideas for reviewers to make decisions. To further guide the reviewing process such that the work aligns with community standards, notes are summarized into broad topics and structured according to community standards, which includes aspects of summary, strengths, and weaknesses. Given these high-level outlines, reviewers can add their own points or delete any points. If reviewers would like to expand the high-level outline into a detailed outline with complete bullet points, reviewers can click the expand button that synthesized notes into a structured outline, as shown in Figure 5.
4.1.4. DG4: Fact-checking between outline and source notes and self-reflections
When the research team introduced the idea of using AI to facilitate the reading and writing process, participants expressed the need for fact-checking and providing explanations. To avoid biasing the decision to either accept or reject the paper, but encourage justifications, ReviewFlow provided the feature that for each synthesized outline bullet, when the reviewers hover on it, corresponding notes used to generate the topic will be highlighted in the middle column. The pdf on the left column will also auto-scroll to shows the location of the notes, as shown in Figure 6.
Figure 6. When the participants hover on each outline element, ReviewFlow presents the visual mapping between the summarized outline and source notes. This visual mapping facilitate fact-checking and comparison between summarised topics and original notes.
To help participants check whether the tone of the review is appropriate as well as the structure of the written review meet the community standard, ReviewFlow provided self-reflection features right before the participants submitted their review. As shown in the right column of the Figure 5, before participants submit their final reviews, they need to self-reflect on each aspect listed on the reflection card. Participants can go back and revise their draft if they are not satisfied with the review.
4.2. System implementation
ReviewFlow’s front-end is a React web application responsible for displaying the PDF, managing the annotation data, and displaying the text input used to write the review draft. The back-end uses a Flask server which handles the GPT-4 API endpoints, MongoDB endpoints for storing data, and GROBID for PDF data extraction. This data includes all of the parsed text of the PDF, its PDF coordinates, and citations linked to their references.
4.2.1. Contextual cues
After the user uploads a paper PDF on the front-end, a request is sent to ReviewFlow’s back-end server where GROBID is used to parse and extract the paper’s content, creating an XML/TEI structured document with the coordinates and content of the section titles, text body sentences, and inline citations. This extracted data is then sent to the front-end client where ReviewFlow combines GROBID’s section data with PDF.js’s section data to make GPT-4 API calls to generate the contextual reflection questions for each of the paper’s sections.
To generate phase-level cues, when the user highlights text on the PDF and clicks on the button with the light bulb icon, a pop-up appears with the 4 critical review aspects: importance, novelty, validity, and clarity. When the user selects one of the aspects and clicks “Get a Question”, GPT-4 will use the selected aspect, highlighted phases, paragraph of the highlighted phases, and paper abstract to generate cues.
4.2.2. In-situ citation recommendation
When the PDF is initialized, ReviewFlow not only utilizes GROBID’s XML/TEI data to create markers for the section-specific reflection questions but also uses the data to create a citation layer that overlays all in-line citations. When the user hover over each citation, a popup shows the paper title, publication date, DOI link and the tldr description that is searched using Semantic Scholar API (sem, 2015). To recommend corresponding papers, the system calls the Semantic Scholar Recommendation API [^2] https://api.semanticscholar.org/api-docs/recommendations and uses the keywords of the paper together with the venue to retrieve the most similar paper. After checking the retrieved papers are not included in the current paper, the top three papers are added as a citation pop-up.
4.2.3. Notes-to-outline Synthesis
After the user creates multiple notes, a “Summarize Notes” button will appear at the bottom of the notes panel. Upon clicking this button, the front-end client will organize all the user’s notes data including the corresponding highlighted paper content, selected tags, note text, and paper abstract. Subsequently, the system server will use this data to prompt the GPT-4 model. We used the few-shot learning paradigm with the corresponding prompts (Liu et al., 2023b): ”Please create three important topics on the paper’s [strengths] and [weaknesses] with less than ten words that combine and summarize the user notes.” This output is then formatted in JSON and streamed directly into the front-end’s text-editable draft panel with a summary and topics organized on strengths and weaknesses. When the user clicks on the ”Expand” button at the bottom of the draft panel, all of the text in the draft text input box, notes, and paper abstract is sent to the back-end server where a GPT-4 API call is requested. The GPT-4 API call will then respond with a streamed JSON-formatted output with more details based on the user’s notes and topics specifically for the Strength and Weakness sections.
5. Method
We designed a within-subjects experiment to answer the questions below:
  • RQ1: How does ReviewFlow affect participants’ final written review quality, compared to the same system with no intelligent scaffolding?
  • RQ2: How does ReviewFlow affect participants’ workflow, in terms of time and engagement?
  • RQ3: What benefits and challenges do participants perceive with ReviewFlow’s intelligent scaffolding?
5.1. Study Design
We conducted a within-subjects experiment with 16 participants where each participant will experience both conditions in two sessions separately. We counterbalanced the order of two conditions and papers using Latin Square to minimize the potential order effect. To reduce knowledge transfer and minimize fatigue, we scheduled the two study sessions of each participant at least 24 hours apart.
5.1.1. Participants
Before the study, we sent out a pre-study survey to participants to collect information about their expertise, research topics, review experience, ML knowledge level, and demographics (e.g. age, gender, race). We recruited 16 participants who have experience in conducting academic research for at least two years and have zero to two years of conference peer review experience. We advertised recruitment messages to colleagues, mailing lists, communication channels, and social media and recruited participants from four universities across the US. Using a snowball sampling approach, we asked participants to refer their friends and colleagues. Participants have on average 1.2 years of writing and submitting academic papers. All participants were compensated for \$20 per hour. The study is IRB-approved.
5.1.2. Procedure
Before the user study, participants filled out a pre-survey that captured their previous experience in reviewing and reviewing papers on ML conferences and their expertise in ML to join the study through an online teleconferencing tool. We asked participants whether they had read the paper before to make sure they all were seeing the work for the first time. In the beginning, participants read and virtually signed our IRB-approved consent form. Then they conducted reviews on two papers using different versions of our interface (ReviewFlow or Baseline). To counterbalance the order effect, we randomized the order of the control condition and the experiment condition for each participant, so half participants encountered ReviewFlow in their first session, and the other half experienced it in their second session. For each session, we followed the review process used at the CSCW conference, where we provided the paper draft and all three individual reviews with their ratings and confidence scores.
For both sessions, participants were told to spend around 60 minutes – and no less than 30 minutes – on the review, but they could take as much time as needed. In the pilot study with 4 participants who had reviewing experience before the study, we found that participants can finish writing reviewing within 45 minutes. Before each condition session, we gave the participants a quick 2-minute demo of the interface, while in the experiment session demo, we also introduced the functionality of ReviewFlow. After two sessions, the research team asked the participants to fill out a post-survey to evaluate the system, such as their satisfaction with each feature and their trust in the system. The research team then conducted a 15-minute semi-structured post-interview to ask open-ended questions about their experience, perceptions, and feedback. The post-interviews were video recorded with participants’ permission and were transcribed into text for later analysis. Each session takes around 75 minutes.
5.1.3. Paper Selection Process
We collected recent one-year papers from HCI-related conferences including CHI, CSCW, UIST, UBICOMP, IUI, DIS using the list provided by  [^3] http://www.conferenceranks.com/. To make our study time short so that people have the energy to finish the task, we filtered papers that had fewer than 3000 words. Then we filtered out papers that have technical terms and jargon. The two papers we selected need to have similar lengths, similar difficulties and from the same conference. Combining all the criteria above, we selected two comparable papers from CSCW Companion that were used for the within-subject experiment. We measured participants’ existing knowledge of the two papers as a control variable. During the study, participants described their knowledge of each paper’s topic from not familiar as 1 to very familiar as 7. All participants reported that they had never read or remembered the papers before.
5.1.4. Measures
We collected a mix of quantitative and qualitative data, including each participant’s log data that captured their interactive behaviors with the system, the final review artifacts for each paper (N=32) that recorded the final review, the post-survey (N=32 and the interview transcripts (N=16). The research team analyzed these combined sources of data to reveal insights towards our research questions.
Quality Ratings on the Final Peer Review
To measure the quality of the review, we recruited two experts who have conducted research for more than three years of review experience in HCI/CSCW conferences. They counted the number of strengths and weakness in the review (a proxy for coverage) and rated the quality of all final reviews (N=32) with a five-dimension rubric based on reviewer guidelines for CSCW conference and previous research (Yuan et al., 2021):
  • Tone: The tone of a peer review is always encouraging and respectful. A good review also highlights what authors have done well.
  • Comprehensive: A good review is always well-organized in structure, which includes a summary, strengths, weaknesses, and a clear description of concerns.
  • Constructive: A good review usually provides constructive suggestions. Following the weakness, reviewers usually will provide actionable items that the author can work on to improve the paper’s quality.
  • Justified: A good review justifies specific reasons for their decisions. Avoid providing a decision without any supporting evidence.
Two experts also rated these five dimensions on a simple seven-point scale (1-7). Each expert first read one paper and wrote a peer review of the paper. After this process, the research team first provided two examples and provided instructions for them to rate and discuss until they reached a consensus on ratings based on the instructions. Then they rate each dimension of the review on the paper they wrote independently (N=16). The inter-rater reliability between two experts on all 32 data is moderate where Krippendorff’s alpha are higher than 0.50 (Krippendorff, 2018). The research team then used the average scores by two experts for each paper dimension.
User Interaction Data
To measure participants’ interaction with the tool, we instrumented the interface to collect a range of user activity log data. We collected two timing measures – how long each participant took to finish the review session and how long each participant spent editing the review within the text box. The entire review session time includes reading time and writing time. The ReviewFlow interface also collected interaction data to indicate how much each participant interacted with each scaffolding feature, including how many times they answered the contextual cues, clicked on the citation summary, checked the citation pop-ups and used the outline summarization feature.
User Preferences and Reactions
To evaluate users’ preferences for the ReviewFlow experience compared to a plain review Editor, we asked participants to fill out a short post-study survey. The survey asked participants to directly compare the perceived usefulness, enjoyment, easiness, and sense of control between the ReviewFlow and baseline system. Then we specially asked participants to evaluate each of the machine-generated features. The survey collected their 7-point Likert scale ratings on perceived usefulness, and perceived accuracy for each feature. We collected the level of cognitive demand using NASA TLX on a scale of 1-7 (Hart and Staveland, 1988). We further collected the feeling of control and collaboration on a scale of 1-7.
Previous research showed that scaffolding can promote learners’ self-efficacy (Yantraprakorn et al., 2013). Here, we further measure whether using the scaffold system can improve novice reviewers’ self-efficacy and knowledge. We asked each participant to report self-efficacy after using the ReviewFlow system and the Baseline system by answering the question “How confident are you in your ability to write a conference peer review next time after using the system( 1 as not confident and 7 as very confident)”.
After the post-survey, we conducted a 15-minute semi-structured interview with all participants to capture their overall thoughts as well as specific perceptions of machine-generated highlights and summaries. For example, the research team asked “What do you think of the difference between the task with and without the support of ReviewFlow”, and “Which function did you use the most? Which one did you like the most?”, and “What concerns did you have when using ReviewFlow?”.
5.2. Analysis
Quantitative data analysis
To measure the effect of the ReviewFlow system on each dimension of the review quality (eg. summative, coverage, justified), we conducted repeated measure ANCOVA tests. We used paper id, the order of experiment conditions (whether ReviewFlow was used for the participant’s first or second paper), the knowledge level of each paper topic, and the length of each review as co-variants.
To measure the effects of the experiment condition on the time they spent reading independent reviews and writing reviews taking the consideration of the differences between the paper topic and order effects, we ran a repeated measure ANCOVA using the paper id, the order of experiment condition and the knowledge level of each paper topic as co-variants.
Qualitative data analysis
All semi-structured interviews with participants are recorded and transcribed. Two researchers conducted iterative open coding on the transcripts using Dovetail [^4] https://dovetailapp.com/ and conducted thematic analysis on the transcripts (Braun and Clarke, 2006). They open-coded the data by identifying topics mentioned by the participants. Initial codes were combined into preliminary themes, which were discussed among the research team. After iteratively discussing the code themes, researchers derived the final themes around: participants’ reactions to each feature, their overall perceptions of the ReviewFlow system, and their concerns about using the system.
6. Results
We report our results from the within-subjects experiment and the interview study. In the within-subject experiments, across both conditions, participants spent an average of 38 minutes writing a review with an average length of 243 words written. 59.3% of participants chose to accept decisions, which is consistent with the original decisions for the two papers. Our findings suggested that ReviewFlow provides more guidance to novice reviewers and can make the review process more useful and engaging.
6.1. RQ1: ReviewFlow helps improve the review draft quality
6.1.1. ReviewFlow helps participants write more comprehensive reviews, as rated by experts
To compare the quality of written review in both conditions, we performed repeated measures ANCOVA to examine the effect of the two conditions on the review length as well as each quality measure. We control the order of experiment condition and the knowledge level of each paper topic as co-variates. As shown in Figure 7, we found that the written reviews in the ReviewFlow condition are statistically significantly more comprehensive reviews than the baseline condition(F = 3.55, p = 0.04 *** ). We found no significant interaction effect between the order of the experiment conditions and no statistically significant differences between the knowledge of the two papers.
Figure 7. Experts evaluation on the quality of review written by ReviewFlow versus Baseline.
We also found that participants wrote longer reviews and called out more strengths and weaknesses in the paper in the ReviewFlow condition. However, there is no significant improvement in the constructiveness of the review. This indicates that the ReviewFlow system can scaffold participants to capture more pros and cons but still cannot help participants write constructive items towards each weakness. In addition, we asked the participants to self-rate their satisfaction with the review quality. Participants are more satisfied with the reviews written in the ReviewFlow condition.
ReviewFlow Baseline
Proxy of review quality
Num of strengths
2.38(0.19) 1.91 (0.17)
Num of weaknesses
2.62(0.22)) 2.08(0.27)
Length(words)
250.8(19.4) 235.8(23.5)
Self-rated satisfaction on the review
4.35(0.4) 4.0(0.34)
Time-on-task
reading time (minutes)
26.9 (1.9) 19.2(2.1)
writing time (minutes)
15.5(2.2) 15.7(2.4)
Table 1. Proxy of written review quality includes number of strengths, number of weaknesses, and length of the review. Participants self-perceived review quality. Time on tasks in both conditions
6.2. RQ2: Longer interaction duration with ReviewFlow but improves users’ self-efficacy
6.2.1. Participants took a longer time to read papers with ReviewFlow but a shorter time on drafting the final review
We performed repeated measures of ANCOVA to examine the effect of the two conditions on writing time and control the length of the review as a co-variate. The results did not show a significant difference. We observed that ReviewFlow reduced review writing time but the reading time is higher compared to the baseline, as shown in Table 1. However, even though participants spent more time on reading, they reported that “ReviewFlow saved me more time on writing review”. One participant reflected on the reasons that “I feel actively receiving these pop-up questions make me spend more time on reading the paper, judging it from multiple dimensions and taking notes, but I think it did save me time on the review drafting since I don’t need to go through all the notes again”(P2).
6.2.2. Participants reported to have higher self-efficacy after experiencing ReviewFlow than the Baseline
To compare self-efficact in both condition, we conducted the Mann- Whitney U Test between each participant’s ratings on self-efficacy given the sample is small (McKnight and Najab, 2010). Results showed that self-efficacy ratings in the ReviewFlow (M = 4.92, STD = 0.40) is significantly higher than the Baseline (M = 3.92, STD= 0.36, p = 0.05 ***)
6.2.3. Participants’ took more notes in the ReviewFlow and used the features actively
Participants took significantly more notes in the ReviewFlow system (M=12.9, STD = 3.7) than baseline environment (M = 9.1, STD = 3.9, Wilcoxon signed-rank test, Z= 103.0, p ¡= 0.01, ***** *). This indicates that participants are more engaged in the reading process with ReviewFlow and not only reading text but also critically reflecting on the content. All participants used the notes-to-outline features which helped them summarize notes into an outline. Three participants did not expand the high-level outline to a detailed outline.
# of participants Avg frequency useful accuracy
Used the section-specific reflection question 14 3.8 5.4 4.9
Used the contextual reflection question on highlight 10 1.6 3.8 4.3
Checked the in-situ citation support 10 1.0 4.6 6.2
Checked the missing citation support 9 1.0 2.5 3.4
Used the summarization notes to high-level outline 16 1.12 6.3 5.2
Expanded the high-level outline to detailed outline 13 1 5.2 4.3
Table 2. Usage of scaffolding features and participants’ rating on the usefulness and accuracy. Average frequency is the average number of times the feature was used. The notes-to-outline feature is used by all participants, while some participants did not expand the high-level outline to a detailed outline.
6.3. RQ3: Participants perceived ReviewFlow as useful in the review workflow
After participants had experienced the two systems, we asked participants to compare them overall directly. As shown in Figure 8, participants highly prefer ReviewFlow over the baseline environment. 92.9% of participants perceived the interface as enjoyable to use. Participants mentioned that the scaffolding features make the review process more engaging and less boring (N=5).
Figure 8. Participants’ reactions to the two systems in terms of easiness, enjoyment, and efficiency
Figure 9. Participants’ perceptions of the engagement, cognitive workload and feeling of control in two conditions
We further measured participants’ perceptions of conducting the task together with the system. In the post-survey, we asked participants to rate cognitive workload, such as distraction and engagement, the feeling of control and collaboration with the interface, and perceived learning outcomes with the interface, as shown in Figure 9. The add-on scaffolded features bring in some distraction since 57.1% of the participants believe that the interaction with the interface distracts them from the review. Interestingly, participants reflected that “It is actually not a bad distraction, since you don’t want to keep reading the paper and get nothing out of it. This distraction is like a pop-up that keeps telling you about review criteria from different places”(P8). Even though the current system uses AI-generated features, such as notes-to-outline synthesis, all participants agree that they still have the control over writing process. One participant mentioned the reason is that “It just synthesizes what I have already written, and if I want, I can just delete them”(P9).
64.3% of participants have the feeling of collaborating with the interface. Participants identified the collaboration mainly happened in the process of answering reflections on each section by asking and answering questions and note-to-outline synthesis. The two-step process where participants can first summarize notes into a high-level outline and then edit and expand it provides participants the feeling of “iterating with an assistant”(P8). When we asked participants their reason for being biased by the outputs, most participants thought that the notes synthesis did not bias them in the current session, but they worried that “lazy people might randomly create notes and use the outline to make a decision ”(P6). In addition, all participants believe that the ReviewFlow interface helped them learn the peer review process.
6.3.1. Participants have different preferences on the scaffolding strategies
As shown in Table 2, participants perceived that section-specific reflections were more useful and more accurate than the contextual question on the highlighted text. Participants reflected (N=2) that the uncertainty on the contextual questions is high since the questions’ quality depends on how much text they highlight. P15 also reflected that “I kind of have a question in my mind when I highlight certain parts of the text, so if the question does not match with the question I was thinking, I feel this is not useful”. We also found that many people did not pay attention to the missing citation support, since they did not focus on evaluating the related work section while reading. The most useful feature perceived by participants is the notes-to-outline synthesis feature where participants reflected that it is “very easy to get a sense on top of my notes”(P3). Participants reflected that the detailed outline bullets contain the same meaning as the high-level outline topic which makes it less useful. We further conducted an in-depth analysis of how participants used and perceived different types of scaffolding strategies:
Contextual reflections cues worked as an expert to guide reviews evaluated from different criteria
Participants described that their process of using the section-specific reflections is slightly different from the process of using context-specific reflections. P11 described that “I first check out the questions real quick before each section, and then I dive into the reading for this section. After I finish reading it, I come back and revisit these questions and reflect on how I feel”. While, for contextual questions, “it is more like I feel confused or not sure what I should think about, I will highlight and use those questions to guide me”(P11). P14 described the section-specific question as “an expert who is sitting there that keeps giving you guided questions and prompts you to think”(P14).
Notes-to-outline synthesis helped participants structure their notes according to community practices
All participants used the notes-to-outline synthesis. Participants reflected that they prefer to have the notes in a “structured version” and that they don’t need to conduct the mapping by hand (N=5). However, there are three participants did not use the expand button to obtain the detailed outline. Two participants intentionally avoided using it since they were worried the detailed bullets generated by AI may mislead them in their decision-making.
6.3.2. Participants’ potential concerns of using ReviewFlow
When we asked participants about their main concerns about using ReviewFlow, they mostly concentrated on the potential errors that AI can make in the notes synthesis process. Participants did not express severe concerns in the study session since the generated outline is mainly reusing or summarized their notes. However, they were still worried that the nuanced tone in the outline produced by AI may exaggerate the weakness of the paper and influence reviewers’ perceptions. For instance, one participant mentioned that if you wrote the notes as “the details around study participants seems a little bit unclear”, but the AI turns it into “this paper lack clarity”, which may be misleading to the reviewer. Moreover, participants expressed their concerns of other reviewers in the community. They are worried that “lazy reviewers or last minute reviewers may randomly leave notes” and then use the tool to generate an outline. This misuse of the tool will not be fair to the paper authors and the research community in general.
7. Discussion
Advances in large language models are changing how work gets done. Our research explores how we might integrate intelligent scaffolding in a way that informs and guides novices and steers away from biasing the underlying judgement. To inform the design of intelligent scaffolding, we conducted a formative study where we learned what strategies and practices experts adopt, as well as, what challenges novices face when writing peer reviews. Our ReviewFlow system aimed to leverage LLMs to provide timely considerations while reading and assessing a submission and structural support when composing a review. Our within-subject experiment (N=16 participants) found that novice reviewers not only preferred ReviewFlow over the baseline system, but they also wrote longer and more comprehensive reviews, as judged by experts. Using ReviewFlow participants spent more time reading the paper and took more detailed notes. Those notes were aided by ReviewFlow’s contextual cues, and helped participants call out more strengths and weaknesses. Using ReviewFlow, participants were more satisfied with their reviews (according to self-ratings) and attributed this to the timely cues and a useful workflow.
In comparison with general peer review practices (Wang et al., 2020; Shah, 2022), participants using ReviewFlow allocated more time to reflect on each section and evaluating the paper’s quality. Interestingly, 40% of participants opted for rejection decisions in both conditions, contrary to the original decisions for the two papers. A typical peer review not only contains the paper summary and its contribution but also raises weaknesses from different aspects together with constructive feedback or thought-provoking questions (Moore, 2013). However, participants in both conditions still found it hard to provide constructive feedback on each weakness, even with example reviews. Perhaps future systems could incorporate additional features into the intelligence scaffolding, such as presenting examples of improvement suggestions from similar prior reviews or offering feedback to the reviewers after they have an initial draft (Krause et al., 2017).
7.1. What role should machine intelligence play in this review process?
AI has come a long way since the early and annoying attempts at supporting work (e.g. Clippy). With the development of LLMs, AI can automate or augment many aspects of knowledge work, including information discovery, sensemaking, and writing (Palani et al., 2023; Liu et al., 2023a; Zhang et al., 2023). While LLMs have become incredibly valuable, the risk now might the risk now might be the tendency, especially among tech-focused innovators, of taking automation too far and dehumanizing work. Prior research has explored whether models can predict a paper decision or even draft a review (Liang et al., 2023). Our ReviewFlow system was designed with the intention of balancing the use of technology with human values. Our system strikes this balance by taking cues from the learning sciences research on scaffolding (Saye and Brush, 2002; Holton and Clarke, 2006). The goal of ReviewFlow is not to produce reviews, per se, but to convey an understanding of how to think while reading a submission and writing a review. Our values place more emphasis on training and preparing novices (perhaps for their next peer review), not just on getting to a final peer review.
Prior work indicates that well-design scaffolding can help novices operate more like experts (Yuan et al., 2016). The concept of Zone of Proximal Development (ZPD) represents the space between what a learner is capable of doing unsupported and what the learner cannot do even with support (Harland, 2003; Wass et al., 2011). Scaffolding works most effectively when it meets novices in their ZPD, meeting people where they need support. and tapering off when they have internalized the best practices. The scaffolding in ReviewFlow is intelligent and contextual. For instance, users get reflection cues that adapt specifically to the current section, however the scaffolding does not adapt to the user. Future work should focus on making scaffolding more interactive and adaptive to a learner’s current level of knowledge.
ReviewFlow was explicitly designed to support novices, but it does not adapt to someone’s existing knowledge or experience with the task. A longitudinal deployment would provide insight on whether novices continue to prefer ReviewFlow, or whether it is useful for early experiences with peer reviewing. As novices become more proficient, they may eventually prefer to work with less structure (e.g. traditional text editor), although even old timers likely find value in ReviewFlow’s pragmatic support for capturing and synthesizing notes. Future research could extend the current system to be more context-aware and more adaptive to users and then study its use with both novices and experienced reviewers over time to gain design insights for interactive scaffolding. Furthermore, future work could explore other properties of scaffolding, such as the timing of when it’s provided, the communication modality, the format or representations used, or the strategies leveraged, such as comparing generated examples vs. guiding questions.
ReviewFlow incorporates three types of scaffolding to support cognitive activities: contextual scaffolding generates reflection questions to aid in paper reading; knowledge scaffolding focuses on citations to facilitate paper evaluation; and structural scaffolding synthesizes notes into structured outlines to assist in review writing. The current study provided insightsinto user perceptions of different scaffolding by observing their usage of all the features and asking them to rate each feature’s usefulness and accuracy. More rigorous and well-controlled ablation studies could help us understand the underlying factors impacting the ReviewFlow experience. For instance, a study with three experimental conditions, each providing only one type of scaffolding—contextual scaffolding, knowledge scaffolding, and structural scaffolding—would allow for a more comprehensive evaluation of the quality of review writing, task efficiency, and the learning impacts for the user.
7.2. How can intelligent scaffolding support novices across the science ecosystem?
Our study of ReviewFlow indicates novice reviewers can benefit from intelligent scaffolding: it helps them evaluate the submission and write more comprehensive reviews. Participants mentioned how they gradually learned about the research community’s practices and became more confident in their review writing. Beyond just writing reviews, the strategies for intelligent scaffolding built into ReviewFlow have the potential to provide value to other aspects of the peer review ecosystem (Shah et al., 2018). For example, similar intelligent scaffolding can be used to support novice reviewers to revise their papers and write a persuasive rebuttal. The process of meta-reviewing, or summarizing independent reviews, could also benefit from scaffolding since there are new people jumping into that role each cycle.
Future deployments of this type of intelligent scaffolding would require careful consideration and round-table discussions within the research community. Previous research revealed that writing with an opinionated language model can affect participants’ opinions in writing and attitude towards social topics (Jakesch et al., 2023). In our study, even though our scaffolding was carefully designed to allow the user to drive the process and to avoid biasing the decision, we still heard participants who were concerned about other people using such a system. Interestingly, very few participants were concerned that their own decision or writing was being biased by the AI, instead, they worried about how others would appropriate the intelligent support and how it might erode a community’s trust in the peer review system. To deploy the system in the future and build trust with people in the community, we need to make sure that the system is robust and trustworthy (Wang et al., 2023). A real-world deployment would need to have a large-scale consent process and commitment from key stakeholders.
Beyond its application in peer reviewing, intelligent scaffolding could support a range of of complex knowledge work in science, such as paper reading, literature reviewing, sensemaking, and paper writing (Palani et al., 2023; Liu et al., 2023a; Zhang et al., 2023; Lin et al., 2024). Notable projects like the semantic reader project have developed interactive and dynamic reading interfaces to aid paper reading and citation discovery (Lo et al., 2023; Kang et al., 2022; August et al., 2022; Fok et al., 2023). PaperPlain introduced key questions for readers to consider, highlighting corresponding answers within the paper content (August et al., 2022). These studies suggest broader possibilities for incorporating intelligent, process-driven scaffolding into science work.
7.3. Ethical considerations of AI scaffolding for academic review
Generative AI and LLMs introduce numerous ethical considerations in the design of human-AI collaboration systems. In the context of academic publishing and peer review, a primary concern involves the potential violation of academic integrity and harm to authorship when directly incorporating automatically generated content into writing artifacts. We try to mitigate this in ReviewFlow by constraining the use of LLMs to primarily sensemaking activities, like guided note taking. In the final step, ReviewFlow synthesizes an outline, not paragraphs, even though LLMs are quite capable of doing so. To limit data sharing, ReviewFlow only sends the LLM the reviewer’s notes and highlighted segments, not the full paper. While individual reviewers may still choose to (mis)use LLMs in these ways, ReviewFlow subtly prioritizes learning and thinking over getting the job done fast.
Another concern revolves around the tendency of LLMs to create inaccurate information or to misled people(Huang et al., 2019; Jakesch et al., 2023). Inline with our scaffolding strategy, ReviewFlow includes a checklist popup that allows users to engage in self-reflection, with reminders to proofread and fact-check their review. To enhance transparency in the writing process, future systems might explore the development of intelligent highlighting. For example, highlights could directly color-code the portions that were user-generated and those that were automatically generated, providing a clear visual indication and promoting a more transparent and accountable writing process.
Given the uncertainties, future LLM-powered tool designers should consider strategies for inducing users into a “reflective skepticism” around all data, especially those produced by machine models. As exemplified in prior work  (Lai, 2011), this involves fostering a mindset of critical evaluation and thoughtful questioning to counteract the potential limitations, hallucinations, biases, or misinformation that may arise from these language models.
8. Limitations
Our study has several limitations. First, ReviewFlow combined multiple scaffolding strategies into one tool, leaving as future work to understand to what extent each strategy impacted the outcomes. Ablation studies could provide more insight on the effectiveness of each scaffold. Second, we simulated a mock peer review scenario where users had about one hour to read and write a review for a short paper. In practice, as we learned in our preliminary interviews, writing peer reviews can hours. A longer study could give insights on the enduring value of scaffolding. Third, we selected two HCI papers to use in our experiment, but different research domains and communities have different guidelines and standards for reviewing. Deploying the system across research communities may bring new insights into the generalization of scaffolding strategies. Lastly, the underlying machine models will evolve over time, which can impact the performance of ReviewFlow, for better or worse. Continuous updates and adaptations to the latest AI models would be required to maintain the tool’s effectiveness and relevance.
9. Conclusion
This research explore techniques for integrating LLMs into intelligent scaffolding for academic peer reviews. Our formative studies found that expert reviewers adopted a workflow of annotating, note-taking, and synthesizing notes before writing, while novices lacked perspective on the prior work in the domain and reviewing standards. Modeling the expert workflow, we developed ReviewFlow –LLM-supported workflow that scaffolds novices using contextual reflection cues, in-situ knowledge support, and notes-to-outline synthesis. In a within-subject experiment with 16 novice reviewers, we found that ReviewFlow led to more comprehensive peer review and higher self-efficacy on the task. We further discuss the implication of intelligent scaffolding in knowledge work.